A particle moves in the xy-plane with a velocity vector given by v =3i^+3tj^, where v is in m/s and time t is in s. At t=0 the particle is at the origin. Find the x- and y-coordinates (in m) of the particle at t=2 s.
1
Expert's answer
2020-12-28T15:04:54-0500
"x=x_0+v_xt=0+3(2)=6\ m\\y=y_0+0.5a_yt^2=0+0.5(3)(2)^2=6\ m"